please explain how to solve this linear system by graphing: x+y-5=0
x=5-y

Answer by rfer(16322)

put both in y int form
y=-x+5
y=-x+5
They are the same so the lines are on top of each other.
y int=-5
(0,-5)
(1,-6)
(2,-7)
and so on.